# Can AR-Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision Making? A Comparative Study of AR See-Through vs. 2D Minimap

### v1

@@ 2.4 Hypothesis
According to dual-process theory, human decision-making can be categorized into two distinct cognitive systems: rapid, heuristic-driven Type 1 thinking, and slower, analytical Type 2 thinking [12]. A widely accepted distinction between these systems is that Type 2 thinking demands significantly more cognitive resources than Type 1 [12].
@@
Under conditions of high time pressure, users often lack sufficient cognitive bandwidth to analytically evaluate AI-generated suggestions, potentially resulting in inappropriate reliance – either rejecting correct suggestions or accepting suboptimal ones [48]. This risk is especially pronounced with 2D minimap visualizations, as these interfaces require users to mentally map the information to their corresponding physical environment, imposing significant cognitive load due to reference-frame switching [30]. In contrast, embedded visualizations can significantly reduce or eliminate these spatial mapping demands by directly fusing information within the physical environment, thus potentially fostering more appropriate reliance on AI suggestions.
@@
    AR see-through promotes more appropriate reliance on AI.
@@
    AR see-through improves overall task performance.
@@
    Response times is comparable across user interfaces.
3 Study Design
We conducted a 2 $\times$ 2 within-subjects study to investigate whether AR see-through can foster more appropriate reliance on AI in spatial decision-making. The two independent variables were the type of visualizations – AR (AR see-through) vs. Minimap (minimap) – and the availability of AI assistance (NoAI vs. AI). The study design is detailed below, with fixed values based on a 13-participant pilot study.
3.1 Task Description
We designed a realistic indoor building scenario, where participants acted as busy employees. The goal was to simulate a time-critical spatial decision-making scenario that closely resembles everyday scenarios to the participants. In each trial, participants were tasked with selecting one coffee machine from four available options within a two-floor building within 20 seconds. The objective was to choose the machine that would allow them to obtain a cup of coffee in the shortest possible time. Optimal decision-making in this task depended on two key factors:
  • Walking Distance — the spatial location of each coffee machine relative to the participant’s current position.
  • Queue length — the number of people currently waiting in line at each machine.
Participants accessed information about both the machine locations and queue lengths via one of two visualizations, and in some trials, they also received AI suggestions.
3.2 Spatial Arrangement
  • Floor Relationship – same or cross: Targets were either located on the same floor as the participant or on a different floor (cross). This factor has a strong impact on Minimap conditions as cross-floor targets made estimating walking distances more challenging.
  • Proximity – close or far: Targets were categorized as either close or far in direct proximity (instead of walking distance) to the participant’s location. This factor has an impact on AR conditions, as distant targets were harder to perceive.
The combination of these two factors resulted in four difficulty levels, ranging from Same+Close (easiest) to Cross+Far (most difficult). To maintain perceptual consistency, all four targets within a trial shared the same floor and proximity level; therefore, no one target is perceptually significantly different from others. The participant’s location was then positioned at the center of a circle passing through all four targets.  Figure 2 illustrates an example of the four difficulty levels.
Grouping Arrangements by Participant Location. Ideally, both participant and target locations would vary each trial. However, moving participants between locations would introduce significant overhead and fatigue. To address this, we grouped arrangements by participant location. Specifically, we selected eight participant locations evenly distributed in the building (Figure 1). For each location, we designed four spatial arrangements, each representing one difficulty level, resulting in $4\times 8=32$ unique spatial arrangements. As a result, the arrangements were naturally grouped into sets of four, each set sharing the same participant location but differing in target distributions. Example of one group is shown in  Figure 2.
Target Optimality. Target optimality was determined by two metrics: walking distance from the participant and its queue lengths (i.e., number of people waiting). For simplification, we estimate the total time to obtain coffee from a target using this formula:
$\text{Time}_{\text{coffee}}=\frac{\text{Distance}_{\text{walk}}}{\text{Speed}_{\text{walk}}}+\text{Number}_{\text{people}}\times\text{Time}_{\text{wait}}$
Based on the pilot studies, we empirically set the walking speed at 1m/s and waiting time per person in line at 15s. These parameters were explicitly communicated to participants in the study.
Since the spatial arrangement of targets is fixed, we manipulated the queue lengths to clearly differentiate target optimality. Based on prior research [74], we intentionally included suboptimal choices to effectively evaluate potential participant over-reliance on AI suggestions. Each trial thus included one optimal, two suboptimal, and one worst choice, each differing by at least 10-second $\text{Time}_{\text{coffee}}$. Detailed calculations are provided in the supplementary materials.
3.3 Conditions
With a within-subjects design, each participant completing trials under four conditions: Minimap+NoAI, Minimap+AI, AR+NoAI, and AR+AI. The order of conditions was counterbalanced across participants using a Latin square design to minimize order effects and learning biases. Our primary focus was on participants’ behavior in the AI conditions, while the NoAI conditions served as baselines to better understand user decision-making patterns.
Implementation and Apparatus. Following similar study [10], a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 3.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR see-through visualization was developed using ARKit  [^2] https://developer.apple.com/augmented-reality/ to scan the building and construct a 3D digital twin in Unity  [^3] https://unity.com/, with ARAnchor Manager ensuring proper registration of the digital model with the physical environment. The system is available at our public repository http://public-when-ready.
All trials were completed using an AVP  [^4] https://www.apple.com/apple-vision-pro/ to eliminate device-related differences and provided a high-fidelity immersive experience. Participants were required to have normal vision, use contact lenses, or can use the prescription inserts provided by us.
Figure 1: The 8 participant locations in the study, distributed across two floors of the building. Participants were instructed to complete 4 trials at each location, following the order indicated by the numbered labels. The letter “L” is an abbreviation for location.
Figure 2: An example spatial arrangement for a single participant location, illustrating four distinct difficulty levels: same+close, same+far, cross-floor+close, and cross+far. The participant’s location is marked in red, and colored stars indicate target locations in each difficulty level.
3.4 Participants
We recruited 32 participants (16 male, 16 female; ages 18–28) from a university via email, word of mouth, and posters. The study lasted approximately one hour, and participants received a \$15 Amazon gift card as base compensation. To incentivize accurate performance, participants earned an additional \$0.20 for each trial in which they selected the optimal target, with the bonus capped at \$20.
To minimize bias related to familiarity with the experimental setting, we pre-screened participants to ensure limited exposure to the building used in the study. Specifically, 19 participants (9 male, 10 female) had never visited the building, 12 participants (6 male, 6 female) had been in the building between one and five times, and 1 participant (male) had visited the building between five and ten times.
3.5 Procedures and Design
@@
Figure 3: One example of Latin square-randomized orders across the four conditions. The order of visualizations (Minimap, AR) and AI (AI, NoAI) conditions would be counterbalanced. Each row represents two groups of spatial arrangements within the same condition, each covering four difficulty levels. Each rectangle represents a single trial. Labels (“opt”, “sub”, “worst”) indicate which target the AI suggests in that trial.
  • Introduction (10mins): Participants first provided informed consent and were explicitly informed about the performance-based incentives. Next, participants received a detailed tutorial that explained the study motivation, the AVP’s calibration and usage, and task procedures.
  • Training Tasks (10mins): Participants completed four training trials – one per condition – to familiarize themselves with the task and visualizations. In these trials, after participants made their selections, the optimal target was explicitly highlighted to illustrate optimal decision-making. Participants were encouraged to ask questions and were informed they could withdraw at any point.
  • Actual Tasks (30mins): Participants completed 32 trials, grouped into eight blocks (one block per standing location, see Figure 3). Within each block, they remained physically stationary and completed four trials (one per spatial difficulty). Each condition was tested across two consecutive participant locations (i.e., eight trials per condition). Breaks were allowed during the study.
  • Post-study Interview (10mins): Participants were interviewed with a list of open-ended questions (attached in supplements).
Detailed Trial Procedure. Each trial proceeded through the following steps: The experimenter escorted the participant to the designated starting point and helped them put on the headset. When ready, the experimenter used a controller to initiate the trial, activating the visualization and a 20-second countdown timer. Participants verbally announced their selection (e.g., “Target A”) as soon as they identified the optimal target. Then, the experimenter stopped the timer, automatically clearing the visualization. Participants rated their confidence in their choice using a 1 (Strongly unconfident) to 7 (Strongly confident) scale. Lastly, participants physically pointed toward the actual spatial location of their chosen target. This process was repeated for each trial.
Figure 4: Classification of reliance as appropriate, under-, and over-reliance based on the alignment between AI suggestions and user decisions.
3.6 Measures
We recorded each session through audio and first-person view video. For every trial, we documented and calculated the following measures based on prior research [74, 66]:
  • Accuracy: We defined accuracy as the average point score across trials in each condition, ranging from 0 to 1. Following previous research [74], participants earned points according to their choice quality: optimal choice (1 point), suboptimal choices (0.5 points), and the worst choice (0 points). Each trial contained exactly one optimal target, two suboptimal targets, and one worst target.
  • AI Reliance: Following previous studies [93], we categorized participant reliance on AI as either appropriate or inappropriate (Figure 4). We classified AI reliance as appropriate when participants either followed correct AI suggestions or overrode incorrect ones with better choices. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting correct suggestions in favor of a worse option.
  • Point Error: This measure indicates whether participants correctly pointed to the actual physical location of their selected target. More pointing errors suggest that the visualization falls short in supporting spatial mapping from the digital to the physical space.
  • Confidence: Participants rated their confidence in their decision on a Likert scale from 1 (Strongly unconfident) to 7 (Strongly confident).
3.7 Using Simulated AI
Similar to prior works [74, 10], we used a simulated AI with an average accuracy of 0.75 to ensure controlled experimentation. In each AI condition, the AI suggested the optimal target in five trials, a suboptimal target in two trials, and the worst target in one trial (Figure 3). This setup yields an average accuracy of 0.75 = $\frac{5\times 1+2\times 0.5+1\times 0}{8}$. The order of these suggestions was randomized and counterbalanced across participants, ensuring that optimal, suboptimal, and worst suggestions were evenly distributed across all trials and participants.
Mitigating effects of trust in AI. Following previous work [3, 10], we took steps to neutralize trust effects so that observed differences in reliance could be attributed primarily to the visualizations. Specifically, we avoided anthropomorphizing the AI [33], consistently referring to it as “the AI”. We also described AI outputs as “suggestions” to counteract perfect automation schemas [19] and emphasize that these suggestions could be incorrect. Furthermore, we do not provide decision feedback for each trial [14], and any information about the AI’s accuracy [95].
4 Results
We report both the quantitative and qualitative results from the study.
4.1 Quantitative Results
@@
Figure 5: Summary of quantitative results across conditions. Metrics include (A) accuracy, (B) task completion time, (C) completion time excluding initial search, and (D) confidence ratings. Comparisons are shown between AI-assisted and non-AI conditions for both Minimap and AR visualizations, with significance levels indicated.
4.1.1 Accuracy
@@
Further analysis suggests that AR+AI under-performance stems from inappropriate reliance on AI. The AI system alone achieved a baseline accuracy of 0.75, while human participants in AR+NoAI achieved $Mdn=0.66$. Surprisingly, participants in the AR+AI condition ($Mdn=0.70$) still performed below the AI baseline, suggesting inappropriate reliance on AI. This aligns with findings in prior empirical studies [3]. In contrast, performance in both Minimap+NoAI ($Mdn=0.88$) and Minimap+AI ($Mdn=0.88$) exceeded the AI baseline, indicating that participants were able to appropriately rely on AI by identifying and overriding poor suggestions.

### v2

@@ 4.8. Hypothesis
According to dual-process theory, human decision-making can be categorized into two distinct cognitive systems: rapid, heuristic-driven Type 1 thinking, and slower, analytical Type 2 thinking (Chaiklin2012ThinkingFA). A widely accepted distinction between these systems is that Type 2 thinking demands significantly more cognitive resources than Type 1 (Chaiklin2012ThinkingFA).
@@
Under conditions of high time pressure, users often lack sufficient cognitive bandwidth to analytically evaluate AI-generated suggestions, potentially resulting in inappropriate reliance, either rejecting correct suggestions or accepting suboptimal ones (Lfstrm2023OnTD). This risk is especially pronounced with 2D minimaps, as these interfaces require users to mentally map the information to their corresponding physical environment, imposing substantial cognitive load due to reference-frame switching (Harris2013UnderstandingST). In contrast, embedded visualizations can substantially reduce or eliminate these spatial mapping demands by directly fusing information within the physical environment, thus potentially fostering more appropriate reliance on AI suggestions.
@@
    AR X-ray improves overall task performance, specifically decision accuracy.
@@
    AR X-ray promotes more appropriate reliance on AI.
@@
    AR X-ray shortens response times in decision-making.
5. Results
5.1. Quantitative Results
Prior to analysis, we excluded trials in which participants failed to make a decision within 20 seconds. After filtering, we retained 248 trials in Minimap+AI, 247 in Minimap+NoAI, 237 in X-ray+AI, and 237 in X-ray+NoAI, out of 1024 (32 participants $\times$ 32 trials). For each measure, trials were aggregated by participant and conditions being analyzed. We first tested for normality using the Shapiro-Wilk test. Results indicated that accuracy, point error, and AI reliance significantly violated the assumption of normality, whereas time and confidence did not. Accordingly, we used repeated-measures ANOVA ($\alpha=.05$) to assess the significance effects on time and confidence; then, post-hoc pairwise comparisons were conducted using Holm-Bonferroni corrections. For accuracy and point error, we used Aligned Rank Transform (ART) ANOVA (Wobbrock2011ART) to test significance and use ART-C (Elkin2021ART_C) for post-hoc contrast testing. To compare AI reliance in the Minimap+AI and X-ray+AI conditions, we used Wilcoxon signed-rank tests.
Figure 7. Summary of quantitative results across conditions. Metrics include (A) decision accuracy, (B) task completion time, (C) completion time excluding initial search, and (D) confidence ratings. Comparisons are shown between AI-assisted and non-AI conditions for both Minimap and X-ray visualizations, with significance levels indicated. The “X” denotes the mean, the horizontal line inside each box indicates the median, the box spans the interquartile range (IQR). Outliers are defined as values beyond the upper cutoff $Q_{3}+1.5\times\mathrm{IQR}$ or the lower cutoff $Q_{1}-1.5\times\mathrm{IQR}$.
@@ 5.1. Quantitative Results
5.1.1. Decision Accuracy
We computed the mean accuracy of target selection for each condition at the participant level by averaging across all valid trials for each participant (Figure 7a). The analysis showed a significant effect between Visualizations on the decision accuracy ($F(1,93)=82.07,p<.001$). Post-hoc analysis further revealed that participants were more accurate in the Minimap+AI condition ($Mdn=0.88,IQR=0.13$) than in X-ray+AI ($Mdn=0.70,IQR=0.13$), with a significant effect ($F(1,93)=6.26,p<.001$). Participants were more accurate in Minimap+NoAI ($Mdn=0.88,IQR=0.16$) than in X-ray+NoAI ($Mdn=0.66,IQR=0.13$), with a significant effect ($F(1,93)=6.56,p<.001$). These results reject our first hypothesis H1.
Further observation suggests that X-ray+AI under-performance may stem from inappropriate reliance on AI. Participants in the X-ray+NoAI condition achieved a median decision accuracy of $Mdn=0.66$, which increased to $Mdn=0.70$ in the X-ray+AI condition. However, this improvement still fell short of the AI standalone baseline accuracy of 0.75, suggesting that participants may have relied on it inappropriately. This aligns with findings in prior empirical studies (2021BansalDoesTheWholeExceedItsParts). In contrast, performance in both Minimap+NoAI ($Mdn=0.88$) and Minimap+AI ($Mdn=0.88$) conditions exceeded the AI baseline, indicating that participants were able to appropriately rely on AI by identifying and overriding poor suggestions.
@@ 5.1.1. Decision Accuracy
Figure 8. Number of trials per block in which participants exhibited each type of AI reliance, including (A) appropriate reliance, (B) under-reliance, and (C) over-reliance. Random-selection baselines (in red) are indicated. The “X” denotes the mean, the horizontal line inside each box indicates the median, the box spans the interquartile range (IQR). Outliers are defined as values beyond the upper cutoff $Q_{3}+1.5\times\mathrm{IQR}$ or the lower cutoff $Q_{1}-1.5\times\mathrm{IQR}$.
@@
5.1.2. AI Reliance
Results from the accuracy analysis indicated that participants relied on AI inappropriately in the X-ray+AI condition. Quantitative analysis on the AI Reliance measure confirms this pattern: we observed a significant difference between Minimap+AI and X-ray+AI, with participants showing less appropriate reliance on AI in X-ray+AI (Figure 8a). This finding again contradicts our second hypothesis H2.
Inappropriate reliance on AI can be either over-reliance or under-reliance. To examine which pattern dominated participants’ interactions with AI in X-ray+AI, we separately analyzed trials of over-reliance and under-reliance. Figure 8b and c depict the average occurrences of over- and under-reliance, respectively, per block (8 trials).
To interpret these numbers, we compared the observed behaviors with a random-selection baseline. Among the 8 trials in each block, the AI suggested 5 optimal, 2 sub-optimal, and 1 worst target. If a participant were to choose a target at random (i.e., with equal probability of $1/4$ for each of the four possible targets), the expected numbers of over- and under-reliance trails per block are 0.75 and 4.25, respectively (see Supplemental Materials for calculation details). Compared to this baseline, participants in the X-ray+AI condition displayed fewer instances of under-reliance ($M_{\text{{X-ray}}}=1.63<M_{\text{random}}=4.25$), but notably more instances of over-reliance ($M_{\text{{X-ray}}}=1.41>M_{\text{random}}=0.75$). In contrast, participants in the Minimap+AI condition exhibited fewer instances of both under-reliance ($M_{\text{{Minimap}}}=0.84<M_{\text{random}}=4.25$) and over-reliance ($M_{\text{{Minimap}}}=0.72<M_{\text{random}}=0.75$). These findings indicate that inappropriate reliance in the X-ray+AI condition was primarily driven by over-reliance, consistent with our earlier accuracy analysis (Sec. 5.1.1).
5.1.3. Response Time
We computed the average response time across conditions at the participant level, and observed significant effects (Figure 7b). The analysis showed a significant effect between Visualizations ($F(1,31)=44.30,p<\text{.001}$) and Visualizations $\times$ AI Conditions interaction ($F(1,31)=5.92,p=\text{.02}$) on the response time. Further post-hoc analysis showed that participants responded faster in Minimap+AI ($M=12.08,SE=0.48$) than X-ray+AI ($M=13.03,SE=0.48$) with a significant effect ($\Delta_{M}=\text{-0.94},F(1,31)=-3.12,p=\text{.01}$), and also in Minimap+NoAI ($M=11.76,SE=0.44$) than X-ray+NoAI ($M=13.77,SE=0.44$) with a significant effect ($\Delta_{M}=\text{-2.00},F(1,31)=-6.28,p<\text{.001}$). These results contradict our third hypothesis H3.
Response Time Excluding Initial Search. We observed that participants in the X-ray conditions spent additional time to visually search the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making models (Simon’s model (Simon1977), the drift–diffusion model (ratcliff2008diffusion), the leaky competing accumulator model (usher2001time)) and empirical studies (reisen2008identifying; russo1994eye) suggest that actual decision-making begins only after sufficient information has been gathered. Thus, similar to prior research (Jakobsen2013VISonDisplay), we distinguished the initial search phase from the subsequent decision-making phase. This motivates us to exclude the initial search overhead in the X-ray conditions to better compare its decision-making phase with the Minimap conditions.
Figure 9. (a) and (b) show the direct and walking distance to the orange target; (a) and (b) also indicate the start and end moment of searching for the blue target.
(a) and (b) show the direct and walking distance to the orange target; (a) and (b) also indicate the start and end moment of searching for the blue target.
Following prior visualization-decision work (Padilla2018DecisionMW), we consider that meaningful decision-making start only after at least two options have been identified. Thus, for each trial in the X-ray conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (Figure 9a) to the frame when the second target appeared (Figure 9b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Yet, we acknowledge that this measure is not a theoretically perfect indicator of true decision time. Nevertheless, it provides a scientifically reasonable approximation that is consistent with definitions and procedures used in prior research (Jakobsen2013VISonDisplay; Padilla2018DecisionMW; reisen2008identifying; russo1994eye). We discuss its limitations and possible refinements in Sec. 6.5.
Two co-authors independently annotated these segments frame by frame from the first-person recordings to ensure inter-rater reliability. Inter-rater reliability was assessed using a two-way mixed effects intraclass correlation coefficient, ICC(3,1), which indicated excellent agreement ($\mathrm{ICC}(3,1)=0.995$, 95% CI [0.992, 0.997]) (koo2016guideline). Encoding consistency was ensured through frame-level accurate marking and the clear visual cues afforded by first-person video recordings. Discrepancies were discussed and resolved through consensus. After subtracting the annotated initial search time from the response time, we observed a different pattern (Figure 7c). The analysis showed that there is a significant effect of Visualizations $\times$ AI Conditions interaction ($F(1,31)=7.70,p=\text{.009}$). Further post-hoc analysis indicated that participants in X-ray+AI ($M=11.23,SE=0.48$) spent less time than those in Minimap+AI ($M=12.08,SE=0.48$) with a significant effect ($\Delta_{M}=\text{-0.85},F(1,31)=-3.27,p=\text{.016}$), indicating that, with AI assistance, decisions were made more quickly with the X-ray visualization. No significant difference was observed between Minimap+NoAI and X-ray+NoAI, suggesting that, without AI assistance, participants spent comparable time making decisions across visualizations. In summary, although raw response times were longer in X-ray conditions, a detailed breakdown suggests that this extra time wasn’t entirely spent on the decision-making. This leads to interesting implications that we will further discuss in Sec. 6.4.
5.1.4. Post-Trial Pointing Error Rate
We analyzed participants’ post-trial pointing error rates across the two visualizations (Minimap vs. X-ray) at the participant level, regardless of AI availability. We found a significant effect ($F(1,93)=40.77,p<.001$), with higher error rates in the Minimap condition ($Mdn=0.08,IQR=0.20$) compared to the X-ray condition ($Mdn=0,IQR=0$). This result indicates that the X-ray visualization significantly improved participants’ spatial mapping ability compared to the Minimap.
5.1.5. Self-Reported Confidence
5.2. Qualitative Results
We conducted semi-structured post-study interviews to capture participants’ strategies, rationales, and motivations behind their decision-making across all conditions. Participants are referred to as P1–P32. We analyzed the interviews using reflexive thematic analysis (braun2019reflecting), identifying themes both inductively (bottom-up) and deductively (top-down). Below, we summarize the key themes that emerged from the analysis.
5.2.1. Diverse Decision Strategies Across Visualizations
We observed that participants adopted different decision-making strategies depending on the visualization.
Both Visualizations: People First and Distance First Strategies. Participants primarily adopted one of two strategies: (1) prioritizing fewer people in line, then considering walking distance; or (2) prioritizing shorter distance, then checking queue length. In the Minimap conditions, participants were almost evenly split between these two strategies, with roughly half employing a “people first” (15/32) and the other half a “distance first” (14/32). In the X-ray condition, more participants adopted a distance-first strategy (15 out of 32), while fewer prioritized queue lengths (11/32). The remaining participants (6/32) reported balancing both factors equally without a dominant criterion.
X-Ray: Unique Heuristics and Motivations. While strategy in the Minimap condition appeared to depend largely on participants’ spatial skills, the X-ray condition introduced distinct perceptual heuristics influencing participants to prioritize distance:
  • Egocentric Spatial Imagery. The egocentric view in our AR X-ray appeared to support participants in vividly imagining physical navigation, particularly for complex pathways involving hallway turns or floor changes. For example, P9 noted, “It’s easier to imagine how I should walk to the coffee machine, especially when targets are on a different floor from me…” Such imagination of navigation in the AR X-ray may have led to weigh distance more heavily in their decisions.
  • Visual Proximity Illusions. Participants reported a misleading proximity effect caused by the X-ray view’s “see-through” nature. Because the visualization allows users to see targets behind walls, some targets appeared directly nearby (Figure 9a) even though the actual walking path to reach them was much longer (Figure 9b). This discrepancy often led users to underestimate the true walking distance, particularly when they cannot easily find the path. As P8 noted, “I think this way is shorter but it’s actually longer.” This difficulty occasionally placed additional cognitive burden on participants and may led them to overweight distance.
  • Occlusion Challenges. Some participants (e.g., P2, P7, P11) reported difficulties estimating queue lengths due to occlusion in the X-ray conditions, prompting reliance on distance estimates instead of people counts.
Figure 10. Example of occlusion in the AR X-ray visualization, where overlapping avatars make it difficult to estimate the number of people in the queue, leading to potential misjudgment of wait time.